find the slope -6x+5y=90
Answer by RAY100(1637)

-6x +5y =90
.
rearranging
.
5y = 6x +90
.
divide both sides by 5
.
y = (6/5) x + 18
.
from std form, y = mx +b,,,,m = slope = (6/5)
.
checking, plot a few points ,,,,(0,18),,,,(-15,0)
(y2-y1) / (x2-x1) = 18 /15 =6/5,, ok